OpenStudy (anonymous):

Please help!! If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the third equation. 2x - 3y + z = -4 2x - 2y + z = -1 x - 2y + 3z = -6

OpenStudy (anonymous):

replace \(x\) everwhere you see it in the top two equations by \(2y-3z-6\)

OpenStudy (anonymous):

ok for the first i got y-5z=8

OpenStudy (anonymous):

first one would simplify to \(2y-5z=8\)

OpenStudy (anonymous):

yeah ok you got the idea

OpenStudy (anonymous):

then once get all those what do i do??

OpenStudy (anonymous):

or is those the final answers?? do you have to combine??

OpenStudy (anonymous):

do it for the next equation just like you did for the first one you get \(2y-5z=11\)

OpenStudy (anonymous):

ok 2nd is 2y-11z-12 and 3rd is -6=-6

OpenStudy (anonymous):

that will give you the new system \[2y-5z=8\] \[2y-5z=11\]

OpenStudy (anonymous):

btw this system has no solution

OpenStudy (anonymous):

THANK YOU! :)